# Paragraph-to-Image Generation with Information-Enriched Diffusion Model

### v1

@@ Abstract
Text-to-image (T2I) models have recently experienced rapid development, achieving astonishing performance in terms of fidelity and textual alignment capabilities. However, given a long paragraph (up to 512 words), these generation models still struggle to achieve strong alignment and are unable to generate images depicting complex scenes. In this paper, we introduce an information-enriched diffusion model for paragraph-to-image generation task, termed ParaDiffusion, which delves into the transference of the extensive semantic comprehension capabilities of large language models to the task of image generation. At its core is using a large language model (e.g., Llama V2) to encode long-form text, followed by fine-tuning with LoRA to align the text-image feature spaces in the generation task. To facilitate the training of long-text semantic alignment, we also curated a high-quality paragraph-image pair dataset, namely ParaImage. This dataset contains a small amount of high-quality, meticulously annotated data, and a large-scale synthetic dataset with long text descriptions being generated using a vision-language model.
Experiments demonstrate that ParaDiffusion outperforms state-of-the-art models (SD XL, DeepFloyd IF) on ViLG-300 and ParaPrompts, achieving up to $15\%$ and $45\%$ human voting rate improvements for visual appeal and text faithfulness, respectively. The code and dataset will be released to foster community research on long-text alignment.
The project website is at: ParaDiffusion.
@@ 1 Introduction
  1. 1.
    For the first time, we explore the long-form text alignment challenge for the image generation task, namely, the paragraph-to-image generation task. A comprehensive solution from both data and algorithmic aspects is provided for the challenging setting.
  2. 2.
@@
  4. 4.
    Experiments demonstrate that ParaDiffusion outperforms state-of-the-art models (SD XL, DeepFloyd IF) on ViLG-300 and ParaPrompts, achieving up to $15\%$ and $45\%$ human voting rate improvements for visual appeal and text faithfulness, respectively.
@@ 2 Related Work
Text-to-Image Models. Text-to-image generation [10] require the model to generate an image corresponding to a given textual description. Recently, diffusion-based methods [27, 26, 4, 41, 7, 8, 2, 13] have demonstrated remarkable performance and found applications in various downstream tasks, including controllable generation [42, 29], controlled editing [9], and perception tasks [39, 38]. Stable Diffusion [27] enhances and accelerates the traditional DDPM [10] by conducting denoising processes in the latent space. Imagen [30] firstly incorporates large frozen language models (i.e., T5--XXL [25]) as text encoders for text-to-image generation tasks, demonstrating their significant performance. DALL-E2 utilizes CLIP [24] as a text encoder and a diffusion model [31] as a decoder to address text-to-image generation tasks. Emu [4], on the other hand, introduces new insight that supervised fine-tuning with small but high-quality, visually appealing images can significantly improve the generation quality. RAPHAEL [40], ERNIE-ViLG [6], and Ediffi [1] approach the task of text-to-image generation from the perspective of an ensemble of expert denoisers, exploring potential gains in performance.
Recently, DeepFloyd [5] and PIXART-$\alpha$ [3] further validate the superior text-image semantic alignment capabilities of large language models over CLIP, as they both employ T5 XXL as a text encoder. However, these models still solely explore short-text textual alignment tasks, restricting the text encoder to within 128 tokens. They do not delve deeply into unlocking the full potential of large language models (LLM) and lack exploration of data methodologies. In contrast to these approaches, we further investigate the gains in rich paragraph-image alignment capabilities offered by the state-of-the-art language model (i.e., Llama V2 [35]). Additionally, we provide an effective solution from a data-centric perspective.
@@ 3 Approach
As discussed earlier, we propose a comprehensive solution at both the data and architecture levels for the paragraph-to-image generation task. Thus, this section is divided into two parts: 1) Algorithm Level (§3.1). We introduce a paragraph-to-image generation architecture with diffusion model. 1) Dataset Level (§3.2). we present a high-quality, textual rich paragraph-to-image pairs dataset, where the corresponding textual descriptions can extend up to $400$ words.
@@ 3.1 Algorithm: ParaDiffusion
3.1.1 Text Paragraph Encoder
Prior works [5, 3] directly utilize frozen encoder-decoder language model (i.e., T5 [25]) for short-text to image generation tasks, restricting the text encoder to within 128 tokens. We present a insight that efficiently fine-tuning a more powerful decoder-only language model can yield stronger performance in long-text alignment (up to 512 tokens). Compared to the encoder-decoder architecture of large language models, recent decoder-only models, such as ChatGPT [22], and Llama V2 [35], have garnered more attention and success. There are some evident advantages to using the decoder-only language models as text encoders: 1) Models like Llama and the GPT series, based on the decoder-only architecture, exhibit stronger semantic understanding and generalization capabilities. 2) The training corpora are also nearly four times larger than that of the encoder-decoder language model, i.e., T5 [25]. However, typically, decoder-only architectures are not adept at feature extraction and mapping tasks. Therefore, we propose Paragraph-Image Alignment Learning with Language Model Adaptation to address this issue, adapting the decoder-only model, i.e., Llama V2 [35] for text-to-image generation tasks.
3.1.2 Paragraph-Image Alignment Learning with Language Model Adaptation
Given a text description $y$, the standard Text to Image Diffusion Models is probabilistic model designed to model conditional distributions of the form $p(z|y)$, where a conditional denoising autoencoder $\epsilon_{\theta}(z_{t},t,y);\,t\in{\{1,\dots,T\}}$ is used to learn the reverse process of a fixed Markov Chain of length $T$. The corresponding objective can be simplified to:
@@ 3.1.2 Paragraph-Image Alignment Learning with Language Model Adaptation
where $\tau_{\theta}$ denotes the text encoder. $\mathcal{E}$ refers to a AE [27] for mapping image to latent feature. Typically, during training, we optimize only the Unet of the diffusion model $\epsilon_{\theta}$, while freezing the $\tau_{\theta}$. This is because we consider the text embedding from $\tau_{\theta}$, i.e., CLIP, to be well-suited for text-image alignment tasks. However, when using an LLM as the text encoder, it is essential to consider whether a frozen LLM is appropriate for this context.
@@
where both $\epsilon_{\theta}$ and $\tau_{\Delta\theta(\Theta)}$ can be jointly optimized for learning better text representation. Compared to direct fine-tuning, this strategy offers two advantages: 1) During the training of paragraph-image alignment, it preserves the powerful semantic understanding capabilities of the LLM, preventing knowledge overfitting to simple text-image semantics. 2) Storage and compute efficient. Requires only limited computational resources and incurs no additional inference costs.
@@ 3.1.3 Training Strategy
Stage 1: Pre-train for Short-text and Image Alignment In this stage, we gathered a dataset of 300 million examples to train the Unet of diffusion model with 1.3 billion parameters. Among these, one billion examples were selected from the 5 billion images of the LAION dataset [32]. Additionally, we constructed an internal dataset containing over 200 million examples. Similar to SDXL [23], the model was trained with progressively increasing resolutions. It started with training at a resolution of 256 and was later adjusted to a resolution of 512. The entire pre-training process took 5 days on 56 A100 GPUs.
@@
Stage 3: High-Quality Alignment Data Finally, we created an extremely small but high-quality dataset, namely ParaImage-Small, to further enhance the performance of model with a small learning rate. This dataset consists of 3 manually selected images, each accompanied by human-annotated long-text descriptions. These descriptions provide detailed information about the objects in the images, including their attributes and spatial relationships.
@@ 3.2 Dataset: ParaImage
3.2.1 Image with Generative Captions
@@
3.2.2 ParaImage-Small: Image with Manual Captions
The generated captions from CogVLM cannot be guaranteed to be $100\%$ accurate; therefore, it is essential to create a small but high-quality dataset with manual annotations (i.e., ParaImage-Small). This dataset consists of $3$k, manually selected by annotators from $650k$ exquisite images of LAION-Aesthetics, adhering to common principles in photography. Then, the $3k$ images were annotated with corresponding long-text descriptions, detailing the objects in the images, the attributes of these objects, and their spatial relationships. The details for Paragraph Description Annotation Rule are as follows:
  1. 1.
@@ 3.2.2 ParaImage-Small: Image with Manual Captions
Table 1: Comparison of Data Statistics. ‘short side’ denotes average length of short side for input image. ‘Ave.’ denotes average number of per caption.
@@
Number Short Side Ave. Words Ave. Nouns
@@ 4.1 Implementation Details
Algorithm. Different from the prior works [27, 26, 4], for the first time, we used a powerful decoder-only large language model (i.e., Llama V2 [35]) as the text encoder. To learn the long-term paragraph-image alignment, we adjusted the length of extracted text tokens to $512$ tokens, enabling precise alignment for very complex and semantically rich prompts. For the diffusion model, we follow SDXL [23] and use a Unet of $1.3$B parameters to learn the latent Gaussian noise. Compared to previous work, our training cost is also significantly lower, with the entire model training process only requiring $56$ V100 GPUs for $8$ days. Following LoRA  [11], we apply LoRA to each layer of Llama V2 [35]. At different stages, we employ varying learning rate strategies. For Stage 1, the learning rate is set to $58e\mbox{--}$ – 5 with cosine annealing decay for a single cycle. For Stage 2, the learning rate is $55e\mbox{--}$ – 5, and for Stage 3, it is $51e\mbox{--}$ – 5.
Datasets and Metrics. We use Zero-shot FID-30K on MS-COCO $256\times 256$ [18] and human evaluation on ViLG-300[6] to assess our algorithm. Additionally, considering that the current test prompts focus on short text-to-image generation, ignoring the evaluation for paragraph-to-image generation, we introduced a new evaluation set of prompts called ParaPrompts, including $400$ long-text descriptions. Figure 5 illustrates the distribution comparison of prompt lengths between ParaPrompts and the previous test set. It is evident that previous prompts testing was mostly concentrated on text alignments within the range of 0-25 words, while our prompts extend to long-text alignments of $100$ words or more. Additionally, we present a new insight that longer text descriptions are more challenging, both in terms of visual appeal and text faithfulness. Relevant discussions and comparisons are provided in the supplementary materials. Figure 6 presents the distribution of prompt content, where we categorized prompts into eight scene categories to achieve a more comprehensive evaluation.
@@ 4.2 Performance Comparisons and Analysis
4.2.1 Fidelity Assessment on COCO Dataset
Zero-shot FID-30K on MS-COCO $256\times 256$ [18] is a universal evaluation method for text-image generation tasks. Table 2 presents relevant comparisons between ParaDiffusion and other existing works. Our ParaDiffusion achieved an FID score of $9.64$, demonstrating similar performance to SD XL [23] and PIXART-$\alpha$ [3]. In comparison, RAPHAEL [40] and DeepFloyd-IF [5] achieved better scores, while utilizing larger models with more parameters. Furthermore, we would like to argue that FID may not be an appropriate metric for image quality evaluation, where a higher score does not necessarily indicate better-generated images. Many studies [3, 5, 23], have demonstrated that, instead, the evaluation by human users is a more authoritative measure.
@@ 4.2.1 Fidelity Assessment on COCO Dataset
4.2.2 ViLG-300 [6]
Following the prior works [3, 6], we also conducted a User Study evaluation from the perspectives of visual appeal and text faithfulness on 300 prompts from ViLG-300 [6]. We only chose the recent models with sota performance for comparative evaluation, as involving human evaluators can be time-consuming. Figure 7 presents the related rating proportions from expert evaluators.
Visual Appeal. Our model significantly outperforms DeepFloyd-IF and SD XL-Refiner in terms of visual appeal, while both of these are considered among the best models of the past year. In comparison to PIXART-$\alpha$[3], our model has achieved competitive performance. We want to emphasize that our model did not specifically focus on Visual Appeal. Additionally, PIXART-$\alpha$[3] and our work are concurrent efforts. Therefore, we believe the performance is acceptable.
@@ 4.2.2 ViLG-300 [6]
Text Faithfulness. Figure 7 (right) shows that our model achieved outstanding performance in Text Faithfulness among the four models, with a voting percentage of $13.7\%$. ‘Tie (Same Performance)’ received a very high voting percentage. This is because ViLG-300 includes many simple prompts, leading to consistently good results and making it challenging to differentiate among them. We provide additional cases in the supplementary material to further analyze this situation.
4.2.3 ParaPrompts-400
Figure 8 presents the results on ParaPrompts-400.
Visual Appeal. Similar to ViLG-300 [6], our ParaDiffusion achieved outstanding results in terms of visual appeal on the ParaPrompts dataset, surpassing SD XL and DeepFloyd-IF and approaching PIXART-$\alpha$ [3]. Compared to the performance on the ViLG-300 [6] dataset, there is a decrease in the voting percentage for the ‘Tie’ category. This indicates that in more challenging long-text scenarios, the performance differences among different models become more pronounced.
Text Faithfulness. As for paragraph-guided image generation setting, ParaDiffusion demonstrates a more pronounced advantage in text faithfulness, reaching $60.3\%$, while other models achieve only around $10\%$. This indicates that ParaDiffusion significantly outperforms other models in aligning long-text aspects such as object, object attributes, and object positional relationships. In the supplementary material, we provide additional visualizations and analyses to support this claim.
Table 3: Ablation Study for LLM Adaptation on ParaPrompts-400. We only evaluate Text Faithfulness for the experiment.
Methods # Trainable Params Win (%) Tie (%) Lose (%)

### v3

@@ Abstract
Text-to-image models have recently experienced rapid development, achieving astonishing performance in terms of fidelity and textual alignment capabilities. However, given a long paragraph (up to 512 words), these generation models still struggle to achieve strong alignment and are unable to generate images depicting complex scenes. In this paper, we introduce an information-enriched diffusion model for paragraph-to-image generation task, termed ParaDiffusion, which delves into the transference of the extensive semantic comprehension capabilities of large language models to the task of image generation. At its core is using a large language model (e.g., Llama V2) to encode long-form text, followed by fine-tuning with LoRA to align the text-image feature spaces in the generation task. To facilitate the training of long-text semantic alignment, we also curated a high-quality paragraph-image pair dataset, namely ParaImage. This dataset contains a small amount of high-quality, meticulously annotated data, and a large-scale synthetic dataset with long text descriptions being generated using a vision-language model.
Experiments demonstrate that ParaDiffusion outperforms state-of-the-art models (SD XL, DeepFloyd IF) on ViLG-300 and ParaPrompts, achieving up to $45\%$ human voting rate improvements for text faithfulness. Code and data can be found at: ParaDiffusion.
@@ 1 Introduction
  • For the first time, we explore the long-form text alignment challenge for the image generation task, namely, the paragraph-to-image generation task. A comprehensive solution from both data and algorithmic aspects is provided.
@@
  • Experiments demonstrate that ParaDiffusion outperforms state-of-the-art models (SD XL, DeepFloyd IF) on ViLG-300 and ParaPrompts, achieving up to $45\%$ human voting rate improvements for text faithfulness.
@@ 2 Related Work
Text-to-Image Models. Text-to-image generation [12] require the model to generate an image corresponding to a given textual description. Recently, diffusion-based methods [34, 33, 4, 49, 9, 10, 2, 15] have demonstrated remarkable performance and found applications in various downstream tasks, including controllable generation [50, 36], controlled editing [11], and perception tasks [45, 46]. Stable Diffusion [34] enhances and accelerates the traditional DDPM [12] by conducting denoising processes in the latent space. Imagen [37] firstly incorporates large frozen language models (i.e., T5--XXL [32]) as text encoders for text-to-image generation tasks, demonstrating their significant performance. DALL-E2 utilizes CLIP [31] as a text encoder and a diffusion model [38] as a decoder to address text-to-image generation tasks. Emu [4], on the other hand, introduces new insight that supervised fine-tuning with small but high-quality, visually appealing images can significantly improve the generation quality. RAPHAEL [47], ERNIE-ViLG [7], and Ediffi [1] approach the task of text-to-image generation from the perspective of an ensemble of expert denoisers, exploring potential gains in performance. Recently, DeepFloyd [5] and PIXART-$\alpha$ [3] further validate the superior text-image semantic alignment capabilities of large language models over CLIP, as they both employ T5 XXL as a text encoder. However, these models still solely explore short-text textual alignment tasks, restricting the text encoder to within 128 tokens. They do not delve deeply into unlocking the full potential of large language models (LLM) and lack exploration of data methodologies. In contrast to these approaches, we further investigate the gains in rich paragraph-image alignment capabilities offered by the state-of-the-art language model (i.e., Llama V2 [42]). Additionally, we provide an effective solution from a data-centric perspective.
@@
LLM-based Layout Generation Currently, there is a significant lack of research combining Large Language Models (LLMs) with text-to-image generation models. Only a few studies [8, 23, 6] have explored the use of LLMs for Layout-to-Image Generation, where these works leverage the LLMs to reason over numerical and spatial concepts in text conditions. LayoutGPT [6] proposed to use LLMs as visual planners by generating layouts from text conditions, and thus collaborate with visual generative models. Similarly, LLM-grounded Diffusion [23] and LLM Blueprint [8] focus on extracting critical components from text prompts using LLMs. Currently, these approaches suffer from two main drawbacks: 1) They require multiple models and hand-crafted strategies, making their pipelines complex. They typically involve using an LLM to generate a layout first, followed by layout-based image generation to produce image. 2) The layout-based image generation models utilize box-conditional image generation models [22], which need to be trained on detection datasets. Therefore, these works can be regarded as extensions of box-guided generation. In contrast, our model focuses directly on text-guided image generation tasks and only requires training with image-text pairs. Additionally, our pipeline is very simple, as it is an end-to-end trainable pipeline.
@@ 3 Approach
As discussed earlier, we propose a comprehensive solution at both the data and architecture levels for the paragraph-to-image generation task. Thus, this section is divided into two parts: 1) Algorithm Level (§3.1). We introduce a paragraph-to-image generation architecture with diffusion model. 2) Dataset Level (§3.2). We present a high-quality, textual rich paragraph-to-image pairs dataset, where the corresponding textual descriptions can extend up to $400$ words.
@@ 3.1 Algorithm: ParaDiffusion
3.1.1 Text Paragraph Encoder.
Prior works [5, 3] directly utilize frozen encoder-decoder language model (i.e., T5 [32]) for short-text to image generation tasks, restricting the text encoder to within 128 tokens. We present an insight that efficiently fine-tuning a more powerful decoder-only language model can yield stronger performance in long-text alignment (up to 512 tokens). Compared to the encoder-decoder architecture of large language models, recent decoder-only models, such as ChatGPT [29], and Llama V2 [42], have garnered more attention and success. There are some evident advantages to using the decoder-only language models as text encoders: 1) Models like Llama and the GPT series, based on the decoder-only architecture, exhibit stronger semantic understanding and generalization capabilities. 2) The training corpora are also nearly four times larger than that of the encoder-decoder language model, i.e., T5 [32]. However, typically, decoder-only architectures are not adept at feature extraction and mapping tasks. Therefore, we propose Paragraph-Image Alignment Learning with Language Model Adaptation to address this issue, adapting the decoder-only model, i.e., Llama V2 [42] for text-to-image generation tasks, as shown in Figure 2.
3.1.2 Paragraph-Image Alignment Learning with Language Model Adaptation.
Given a text description $y$ and image $x$, the standard Text to Image Diffusion Models is probabilistic model designed to model conditional distributions of the form $p(z|y)$, where a conditional denoising autoencoder
$\epsilon_{\theta}(z_{t},t,y);\,t\in{\{1,\dots,T\}}$ is used to learn the reverse process of a fixed Markov Chain of length $T$. The corresponding objective can be simplified to:
@@
where $\tau_{\theta}$ denotes the text encoder and its output is a text embedding ($n\times 1280$), as shown in Figure 2. $\mathcal{E}$ refers to a AE [34] for mapping image to latent feature. Typically, during training, we optimize only the Unet of the diffusion model $\epsilon_{\theta}$, while freezing the $\tau_{\theta}$. This is because we consider the text embedding from $\tau_{\theta}$, i.e., CLIP, to be well-suited for text-image alignment tasks. However, when using an LLM as the text encoder, it is essential to consider whether a frozen LLM is appropriate for this context.
@@
where both $\epsilon_{\theta}$ and $\tau_{\Delta\theta(\Theta)}$ can be jointly optimized for learning better text representation. Compared to direct fine-tuning, this strategy offers two advantages: 1) During the training of paragraph-image alignment, it preserves the powerful semantic understanding capabilities of the LLM, preventing knowledge overfitting to simple text-image semantics [13]. 2) Storage and compute efficient. Requires only limited computational resources and incurs no additional inference costs.
@@
Stage 1: Pre-train for Short-text and Image Alignment. In this stage, we gathered a dataset of 300 million examples to train the Unet of diffusion model with 1.3 billion parameters. Among these, $100$ million examples were selected from the 5 billion images of the LAION dataset [39]. Additionally, we constructed an internal dataset containing over $200$ million examples. Similar to SDXL [30], the model was trained with progressively increasing resolutions. It started with training at a resolution of 256 and was later adjusted to a resolution of 512. The entire pre-training process took 5 days on 56 A100 GPUs.
@@
Stage 3: High-Quality Alignment Data. Finally, we created an extremely small but high-quality dataset, namely ParaImage-Small, to further enhance the performance of model with a small learning rate. This dataset consists of 3k manually selected images, each accompanied by human-annotated long-text descriptions. These descriptions provide detailed information about the objects in the images, including their attributes and spatial relationships.
@@ 3.2 Dataset: ParaImage
3.2.1 Image with Generative Captions.
@@
3.2.2 ParaImage-Small: Image with Manual Captions.
The generated captions from CogVLM cannot be guaranteed to be $100\%$ accurate; therefore, it is essential to create high-quality dataset with manual annotations (i.e., ParaImage-Small). This dataset consists of $3$k, manually selected by annotators from $650k$ high-quality images of LAION-Aesthetics, adhering to common principles in photography. Then, the $3k$ images were annotated with long-text descriptions, detailing the objects in the images, the attributes of these objects, and spatial relationships. The details for Description Annotation Rule are as follows:
@@
Table 1: Comparison of Data Statistics. ‘short side’ denotes average length of short side for input image.
@@
Number Short Side Average Words Average Nouns
@@ 4.1 Implementation Details
Algorithm. Different from the prior works [34, 33, 4], for the first time, we used a powerful decoder-only large language model (i.e., 7B Llama V2 [42]) as the text encoder. To learn the long-term paragraph-image alignment, we adjusted the length of extracted text tokens to $512$ tokens, enabling precise alignment for very complex and semantically rich prompts. For the diffusion model, we follow SDXL [30] and use a Unet of $1.3$B parameters to learn the latent Gaussian noise. Compared to previous work, our training cost is also significantly lower, with the entire model training process only requiring $56$ A100 GPUs for $8$ days. Following LoRA  [13], we apply LoRA to each layer of Llama V2 [42]. At different stages, we employ varying learning rate strategies. For Stage 1, the learning rate is set to $58e\mbox{--}$ – 5 with cosine annealing decay for a single cycle. For Stage 2 and Stage 3, the learning rate is $55e\mbox{--}$ – 5 and $51e\mbox{--}$ – 5, respectively.
Datasets and Metrics. We use Zero-shot FID-30K on MS-COCO $256\times 256$ [24] and human evaluation on ViLG-300[7] to assess our algorithm. ParaPrompts. Additionally, considering that the current test prompts focus on short text-to-image generation, ignoring the evaluation for paragraph-to-image generation, we introduced a new evaluation set of prompts called ParaPrompts, including $400$ long-text descriptions. Figure 5 illustrates the distribution comparison of prompt lengths between ParaPrompts and the previous test set. It is evident that previous prompts testing was mostly concentrated on text alignments within the range of 0-25 words, while our prompts extend to long-text alignments of $100$ words or more. Additionally, we present a new insight that longer text descriptions are more challenging, both in terms of visual appeal and text faithfulness. Relevant discussions and comparisons are provided in the supplementary materials. Figure 6 presents the distribution of prompt content, where we categorized prompts into eight scene categories to achieve a more comprehensive evaluation. Text Faithfulness of Sentence Level. To further evaluate the text faithfulness performance, we propose a more fine-grained metric, where caption of ParaPrompts-400 are divided into $3,945$ sentence fragments. Then, an open-domain grounding model  [17] is employed to segment the related content for each sentence. Finally, two evaluators manually determine the precision (whether the relevant content is present in the synthesized image).
@@ 4.2 Performance Comparisons and Analysis
4.2.1 Fidelity Assessment on COCO Dataset.
Zero-shot FID-30K on MS-COCO $256\times 256$ [24] is a universal evaluation method for text-image generation tasks. Table 2 presents relevant comparisons between ParaDiffusion and other existing works. Our ParaDiffusion achieved an FID score of $9.64$, demonstrating similar performance to PIXART-$\alpha$ [3]. In comparison, RAPHAEL [47] and DeepFloyd-IF [5] achieved better scores, while utilizing larger models with more parameters. Furthermore, we would like to argue that FID may not be an appropriate metric for image quality evaluation, where a higher score does not necessarily indicate better-generated images. Many studies [3, 5, 30], have demonstrated that, instead, the evaluation by human users is a more authoritative measure.
@@
Figure 9: User study for LLM Blueprint [8] and ParaDiffusion on 400 prompts of ParaPrompts. ‘Tie’ indicates that the image quality of the four models appears similar.
Figure 10: More Visualization Comparison for Layout-based models (LLM Blueprint [8] and LLM-grounded Diffusion [23]) and ParaDiffusion. Compared to layout-based methods, our ParaDiffusion achieves better text faithfulness, especially in generating overlapping objects.
Figure 11: Visualization Comparison for LLM Blueprint [8] and ParaDiffusion. The effect of ParaDiffusion is better than LLM Blueprint in terms of both Visual Appeal and Text Faithfulness.
4.2.2 ViLG-300 [7].
Following the prior works [3, 7], we also conducted a User Study evaluation from the perspectives of visual appeal and text faithfulness on 300 prompts from ViLG-300 [7]. We only chose the recent models with sota performance for comparative evaluation, as involving human evaluators can be time-consuming. Figure 7 presents the related rating proportions from two expert evaluators. In cases where there was ambiguity in the voting, the two evaluators discussed and reached a consensus. Visual Appeal. Our model significantly outperforms DeepFloyd-IF and SD XL-Refiner in terms of visual appeal, while both of these are considered among the best models of the past year. In comparison to PIXART-$\alpha$[3], our model has achieved competitive performance. We want to emphasize that our model did not specifically focus on Visual Appeal. Additionally, PIXART-$\alpha$[3] and our work are concurrent efforts. Therefore, we believe the performance is acceptable. Text Faithfulness. Figure 7 (right) shows that our model achieved outstanding performance in Text Faithfulness among the four models, with a voting percentage of $13.7\%$. ‘Tie (Same Performance)’ received a very high voting percentage. This is because ViLG-300 includes many simple prompts, leading to consistently good results and making it challenging to differentiate among them. We provide additional cases in the supplementary material to further analyze this situation.
Table 3: A fine-grained metric for the precision of sentence fragments (Text Faithfulness) using LISA [17] on ParaPrompts-400.
Methods SD XL DeepFloyd-IF PixArt-$\alpha$ ParaDiffusion
Precision $25.8\%$ $27.9\%$ $29.4\%$ $51.3\%$
Table 4: Evaluation for Objects, Attributes, and Locations on ParaPrompts 400. We randomly sample 50 samples from ParaPrompts 400 and conduct a 1-5 scoring assessment, with 5 being the highest score.
Methods Objects Attributes Locations
LLM Blueprint [8] 3.4 3.1 4.2
ParaDiffusion 3.9 4.1 4.0
@@
Figure 13: Prompt workflow for detailed concepts in image descriptions.
4.2.3 ParaPrompts-400.
Figure 8 presents the results on ParaPrompts-400. Visual Appeal. Similar to ViLG-300 [7], our ParaDiffusion achieved outstanding results in terms of visual appeal on the ParaPrompts dataset, surpassing SD XL and DeepFloyd-IF and approaching PIXART-$\alpha$ [3]. Compared to the performance on the ViLG-300 [7] dataset, there is a decrease in the voting percentage for the ‘Tie’ category. This indicates that in more challenging long-text scenarios, the performance differences among different models become more pronounced. Text Faithfulness. As for paragraph-guided image generation setting, ParaDiffusion demonstrates a more pronounced advantage in text faithfulness, reaching $60.3\%$, while other models achieve only around $10\%$. This indicates that ParaDiffusion significantly outperforms other models in aligning long-text aspects such as object, object attributes, and object positional relationships. In the supplementary material, we provide additional visualizations and analyses to support this claim. Additionally, we also provide a more fine-grained metric (refer to Section 4.1), as illustrated in Table 3. It can be observed that even at the sentence level, our model achieves an absolute lead in text faithfulness alignment, reaching $51.3\%$, surpassing the other models by nearly $20\%$.
Figure 14: Visualization Comparison for Detailed Concepts. ParaDiffusion exhibits superior alignment capabilities, effectively managing the complexity of multiple objects and attributes.
Table 5: Ablation of Caption Length for ParaImage on ParaPrompts-400. We prompt CogVLM [44] to generate captions of various lengths. $500$k images of ParaImage-Big are randomly sampled from the original dataset. We only evaluate Text Faithfulness for the experiment. Longer captions improve significantly.
Dataset Size Average Length of Caption ($n$ words)
10 words 50 words 100 words 150 words
ParaImage-B 500k 8.5% 21.1% 29.8% 40.6%
ParaImage-S 3k 16.4% 23.3% 29.1% 31.2%
Table 6: Ablation Study for LLM Adaptation on ParaPrompts-400. We only evaluate Text Faithfulness for the experiment. ‘P.’ refers to ‘Parameters’.
Hyperparameters # Trainable P. Win Tie Lose